OpenStudy (anonymous):

Find Sn for each arithmetic series. a1= 62, d = 12, n = 7 Formula is Sn = n/2(a1 + an) Don't know what the d stands for or how to use it in the formula. Answer is 686.

OpenStudy (callisto):

a7 = a1 + (n-1)(d) = 62 + (7-1)(12) = 134 Sn = (7/2) (62 + 134) =?

OpenStudy (kropot72):

d stands for the common difference between consecutive terms of the progression. In this case the first term is 62 and the second term is 62 +12 = 74

OpenStudy (callisto):

n is the number of term while an is the last term. In this case, the last term is a7

OpenStudy (callisto):

you need d to find the last term an = a1 + (n-1) d